# OpenVLThinker: An Early Exploration to Complex Vision-Language Reasoning via Iterative Self-Improvement

### v1

@@
Abstract: Recent advancements demonstrated by DeepSeek-R1 have shown that complex reasoning abilities in large language models (LLMs), including sophisticated behaviors such as self-verification and self-correction, can be achieved by RL with verifiable rewards and significantly improves model performance on challenging tasks such as AIME. Motivated by these findings, our study investigates whether similar reasoning capabilities can be successfully integrated into large vision-language models (LVLMs) and assesses their impact on challenging multimodal reasoning tasks. We consider an approach that iteratively leverages supervised fine-tuning (SFT) on lightweight training data and Reinforcement Learning (RL) to further improve model generalization. Initially, reasoning capabilities were distilled from pure-text R1 models by generating reasoning steps using high-quality captions of the images sourced from diverse visual datasets. Subsequently, iterative RL training further enhance reasoning skills, with each iteration’s RL-improved model generating refined SFT datasets for the next round. This iterative process yielded OpenVLThinker, a LVLM exhibiting consistently improved reasoning performance on challenging benchmarks such as MathVista, MathVerse, and MathVision, demonstrating the potential of our strategy for robust vision-language reasoning. The code, model and data are held at https://github.com/yihedeng9/OpenVLThinker.
Figure 1: Performance (accuracy) of OpenVLThinker-7B on multi-modal reasoning benchmarks.
@@ 1 Introduction
LLMs have displayed remarkable progress in multi-step reasoning, most notably exemplified by OpenAI’s O1 model [20] and O3-mini. Follow-up studies have sought to replicate and extend these complex chain-of-thought abilities [43, 19, 37, 62, 61], revealing that targeted post-training methods can further improve model performance on challenging tasks. Among these, the recently open-sourced DeepSeek-R1 [13] has shown that reinforcement learning with verifiable rewards is particularly effective at eliciting nuanced self-verification and self-correction reasoning from the LLM, thereby effectively enhancing the overall reliability of mathematical or logical reasoning chains.
While these advances predominantly focus on text-only models, the potential of incorporating similarly complex reasoning strategies within large vision-language models (LVLMs) has remained relatively underexplored. Modern LVLMs merge open-source foundation LLMs [50, 51, 9, 56] with text-aligned image encoders [44, 24], enabling them to process visual contexts and produce outputs upon instructions [26, 10, 11, 7, 27, 2]. These LVLMs have broad vision-language pretraining and strong visual instruction-following ability, but lack complex multi-step reasoning abilities like O1 and R1 models. Meanwhile, explicitly training these models to harness sophisticated CoT reasoning is much less developed than their text-only counterparts [65, 66, 38, 17, 14, 55, 49]. Very recently, concurrent works have begun to investigate how to endow LVLMs with R1-like reasoning capabilities via distillation from text-only reasoning models [5, 67, 57, 18, 23], illustrating that advanced reasoning has great potential to contribute to improved performance on vision-language tasks.
@@
Figure 2: Demonstration of the training process toward OpenVLThinker-7B. We obtain the SFT data for iteration 1 via text-based R1 model that only receives the question and the generated image caption. Subsequently, we apply SFT and GRPO iteratively to leverage new reasoning data from the previous iteration and achieve self-improvement. We also evolve the data sources to progressively include more challenging questions over iterations.
In this paper, we present one of the first explorations aimed at enhancing complex reasoning capabilities in LVLMs and introduce OpenVLThinker-7B, a model trained through iterative self-improvement via alternating SFT and RL. Importantly, we consider the role of SFT primarily as to guide the model in obtaining an initial reasoning structure, while RL serves as the main driver for significant performance gains and improved generalization. We thus employ a warm-start approach, utilizing SFT to establish the initial reasoning framework, rather than directly applying RL from scratch. We distill R1-style reasoning from existing text-only models into the vision-language context using textual image descriptions. This image-to-text conversion inevitably introduces information loss, resulting in a comparatively weaker SFT dataset. Nevertheless, since the primary purpose of SFT in our method is to establish a reasoning structure, we proceed from this adequate starting point. Next, we incrementally refine the model through RL and specifically Group Relative Policy Optimization (GRPO) [47]. The improved reasoning capability is then leveraged to generate the reasoning and answer for the SFT training data in the next iteration that trains the model from scratch (Qwen2.5-VL), creating a loop that enhances the LVLM’s complex reasoning behavior and benchmark performance. We also evolve the data sources to progressively include more challenging questions over iterations, from including mostly simpler questions such as TabMWP [34] to more difficult questions such as Geometry3K [30]. We summarize our method pipeline in Figure 2.
We demonstrate the effectiveness of our approach on challenging multimodal reasoning tasks, including MathVista [33], MathVerse [64] and MathVision [53], which require both precise image grounding and robust multi-step reasoning. Our experiments show that (1) the distilled R1 reasoning structure transfers effectively from text-only to vision-language domains and (2) iterative RL can refine and stabilize these reasoning pathways, ultimately surpassing the baseline performance of standard LVLMs. Beyond highlighting the synergy of SFT and RL for complex reasoning, our findings contribute early evidence toward the promise of bridging R1-style reasoning into multimodal contexts.
{mdframed}
[linewidth=0.5pt]
Assistant:
<think>
Okay, so I have this triangle XYZ with sides labeled as 9x, 4x + 5, and 6x + 3.
The problem is asking me to find the length of side XY, which is 4x + 5.
First, I remember that in a triangle, all sides are equal because it’s an
equilateral triangle. That means all sides are the same length. So, I set the
expressions for the sides equal to each other.
I start with the sides: 9x = 4x + 5 and 9x = 6x + 3.
Let me solve the first equation, 9x = 4x + 5. I subtract 4x from both sides:
9x - 4x = 5, which simplifies to 5x = 5. Then, I divide both sides by 5: x = 1.
Now, I substitute x = 1 into the expression for side XY, which is 4x + 5. So,
4(1) + 5 = 4 + 5 = 9.
I double-check my work. The other sides are 9x and 6x + 3. Substituting x = 1,
9x becomes 9(1) = 9, and 6x + 3 becomes 6(1) + 3 = 9.
Both sides are equal to 9, so the triangle is equilateral with all sides
equal to 9.
</think>
The length of side XY is 9.
<answer>D</answer>
Figure 3: An example reasoning trace of OpenVLThinker. It shows desirable reasoning behavior including self-reflection and self-verification, such as “I double-check my work”. Corresponding image and question are shown in Figure 9.
@@ 2.1 Complex Chain-of-Thought Reasoning
Since the introduction of OpenAI’s O1 model [20], researchers have shown strong interest in reproducing and enhancing the complex reasoning capabilities of LLMs [43, 19, 37, 62, 61], partly due to its superior performance on mathematical benchmarks. Guo et al. [13] introduce the open-source DeepSeek-R1 model and investigate how RL with verifiable rewards can promote advanced chain-of-thought reasoning and reflective behaviors. This development inspired a line of research focused on open-source reproduction [35, 16, 63, 29] and the analysis of such complex reasoning in mathematical problem solving [59, 54, 58, 6]. In parallel, several recent studies have similarly explored the effects of test-time scaling on encouraging more complex model reasoning behaviors [39, 46, 28, 12].
@@ 2.2 Vision-Language Reasoning Model
Recent advancements in large vision-language models (LVLMs) stem from open-source LLMs [50, 51, 9, 56] and text-aligned image encoders [44, 24]. Integrating these components has enabled LVLMs to follow diverse visual instructions and generate meaningful responses [26, 10, 11, 7, 27, 2]. Parallel to the model development, researchers have also been interested in eliciting reasoning chains from LVLMs via prompting [65, 66, 38, 17] or fine-tuning [14, 55, 49].
Concurrent work. Following the introduction and open-sourcing of DeepSeek-R1, there have been several concurrent early efforts aiming to equip LVLMs with R1-style reasoning capabilities [67, 57, 18, 23, 42, 36]. Our work aligns with these early studies, and we contribute our initial results and insights to support and further advance community exploration in this direction. Notably, we propose an iterative self-improving approach toward such multi-modal reasoning.
Figure 4: Iterative performance improvement of our model on MathVista. We note that SFT-Iter(i) is always fine-tuned from the base model Qwen2.5-VL-7B, with its training data generated from GRPO-Iter(i-1). GRPO-Iter(i) is obtained by applying GRPO to train SFT-Iter(i).
@@ 3 Preliminaries
Supervised Fine-Tuning (SFT). We usally use SFT to specialize LLMs for a particular task or domain. This process updates the model parameters $\bm{\theta}$ by providing example responses of desired behavior to the input instructions. Concretely, Given a dataset $\mathcal{D}=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N}$, where $\mathbf{x}^{(i)}$ is the prompt sequence and $\mathbf{y}^{(i)}$ is the desired response sequence. We update $\bm{\theta}$ to maximize the likelihood of producing $\mathbf{y}^{(i)}$ given $\mathbf{x}^{(i)}$. Formally, $\mathcal{L}_{\text{SFT}}(\bm{\theta})=-\sum_{i=1}^{N}\log p_{\bm{\theta}}\bigl{(}\mathbf{y}^{(i)}\,\big{|}\,\mathbf{x}^{(i)}\bigr{)}$. By minimizing the loss, the model learns to produce responses more aligned with the labeled examples.
Reinforcement Learning Fine-tuning (RFT). RL approaches fine-tune LLMs via human preferences modeled under the Bradley-Terry model [40, 8, 47, 1]: $p(\mathbf{y}_{w}\succ\mathbf{y}_{l}\mid\mathbf{x})=\sigma\bigl{(}r(\mathbf{x},\mathbf{y}_{w})-r(\mathbf{x},\mathbf{y}_{l})\bigr{)}$, where $\mathbf{y}_{w}$ and $\mathbf{y}_{l}$ denote preferred and dispreferred responses, respectively, and $\sigma(t)=1/(1+e^{-t})$ is the sigmoid function. The common RL objective under the Bradley-Terry assumption of the reward model $r(\mathbf{x},\mathbf{y})$ is thus
@@
where $\beta>0$ is the KL penalty coefficient. Under this framework, Shao et al. [47] introduced Group Relative Policy Optimization (GRPO) by sampling a group of response trajectories $\{\mathbf{o}_{i}\}_{i=1}^{G}$ from the old policy model $\bm{\theta}_{\mathrm{old}}$ for each query $\mathbf{x}$, with the objective as maximizing:

### v2

# OpenVLThinker: Complex Vision-Language Reasoning via Iterative SFT-RL Cycles

@@
Abstract:
We introduce OpenVLThinker, one of the first open-source large vision–language models (LVLMs) to exhibit sophisticated chain-of-thought reasoning, achieving notable performance gains on challenging visual reasoning tasks. While text-based reasoning models (e.g., Deepseek R1) show promising results in text-only tasks, distilling their reasoning into LVLMs via supervised fine-tuning (SFT) often results in performance degradation due to imprecise visual grounding. Conversely, purely reinforcement learning (RL)-based methods face a large search space, hindering the emergence of reflective behaviors in smaller models (e.g., 7B LVLMs). Surprisingly, alternating between SFT and RL ultimately results in significant performance improvements after a few iterations. Our analysis reveals that the base model rarely exhibits reasoning behaviors initially, but SFT effectively surfaces these latent actions and narrows the RL search space, accelerating the development of reasoning capabilities. Each subsequent RL stage further refines the model’s reasoning skills, producing higher-quality SFT data for continued self-improvement. OpenVLThinker-7B consistently advances performance across six benchmarks demanding mathematical and general reasoning, notably improving MathVista by 3.8%, EMMA by 2.4%, and HallusionBench by 1.6%. Beyond demonstrating the synergy between SFT and RL for complex reasoning tasks, our findings provide early evidence towards achieving R1-style reasoning in multimodal contexts. The code, model and data are held at https://github.com/yihedeng9/OpenVLThinker.
Figure 1: Our OpenVLThinker-7B (in red) performs competitively to large proprietary multimodal models such as GPT-4o and Claude-3.5 (in gray), especially in Math and Perceptron tasks. It outperforms VL base models at the same scale (in blue) and other recently released VL reasoning models (in purple).
@@ 1 Introduction
Proprietary large language models (LLMs), notably OpenAI’s o-series [29] and Google’s Gemini-2.5 Pro [17], have demonstrated impressive multi-step reasoning abilities of planning, reflection, and verification. Recent open-weight models [51, 27, 46, 82, 81] (e.g., DeepSeek-R1 [19] and smaller LLMs like S1 [48] and QwQ-32B [61]) show that reinforcement learning (RL) with verifiable rewards effectively reproduces these advanced capabilities, significantly boosting performance on challenging mathematical and logical tasks.
Unlike text-only LLMs, it remains unclear whether open-source large vision-language models (LVLMs) can effectively adopt similar sophisticated reasoning strategies. Modern LVLMs such as LLaVA-NeXT [37] and Qwen2.5-VL [3] benefit from extensive vision-language pretraining and demonstrate strong visual instruction-following capabilities. However, they rarely demonstrate advanced reasoning behaviors like GPT-o1 or DeepSeek-R1.
Moreover, it is known that reasoning capabilities can generally be distilled from larger LLMs to smaller ones through supervised fine-tuning (SFT) on chain-of-thought demonstrations [35, 31] for text-only tasks. This recipe has been recently applied in distills demonstrations from DeepSeek-R1 (LIMO [76], S1 [48] and OpenThinker [60]) followed by optional RL fine-tuning [77]. However, adapting this method to LVLMs does not work. Proprietary LVLMs, such as OpenAI’s o1/o3 and Google’s Gemini, do not expose their internal reasoning paths, making their outputs unsuitable for distillation. Therefore, most recent attempts are focusing on improving LVLMs through distillation from text-only R1 reasoning models (see discussion in Section 2.2). Unfortunately, our experiments show that naively fine-tuning LVLMs on reasoning paths generated from text-based DeepSeek-R1 with image captions leads to a non-trivial performance drop (see Figure 4), primarily due to a lack of precise visual grounding. Similar observations can be found in [6, 74].
In this paper, we present OpenVLThinker-7B, one of the first open-weight LVLMs that exhibit complex reasoning capabilities in complex vision-language tasks. Specifically, it is trained by iterating between the following two steps:
  1. 1.
    Lightweight SFT. In the first iterations, we distill CoTs using a text-only Deep-Seek R1 given the task question and the corresponding generated image caption. These CoT traces provide demonstrations of reasoning actions, although they do not immediately improve LVLM’s accuracy. For later iterations, we use the LVLM from the previous iteration to produce CoTs on 3,000 data points. This small dataset is sufficient to progressively enhance the model’s reasoning depth.
  2. 2.
    Curriculum RL. In subsequent iterations, we further enhance the LVLM’s reasoning through RL exploration with Group Relative Policy Optimization (GRPO) [54], which splits training into two rounds to form a smooth curriculum.
We found that while the initial step of SFT leads to a performance drop, iteratively alternating between SFT and RL eventually gradually yields a significant performance gain on both reasoning depth and answer accuracy (Figure 2).
Our further analysis shows that the inference-time reasoning behaviors are often triggered by specific tokens (e.g., "wait"). SFT serves as an inductive prior that highlights these reasoning actions. Specifically, it demonstrates the tokens such as “first”, “wait”, “check”, that trigger the model’s planning, reflection, and verification behaviors. Without this SFT step, launching RL from scratch forces the model to search through a prohibitively large space, making reflective behaviors slow to emerge – if they emerge at all. On the other hand, RL plays the critical role in learning the reasoning behaviors, generalizing from training data, and offering a better foundation for the next SFT iteration. The iterative cycle between SFT and RL collaboratively optimizes LVLM’s performance.
@@
Figure 2: Illustration of OpenVLThinker-7B’s training process. We iteratively apply SFT and GRPO to refine the LVLM using reasoning data generated from previous iterations. The data sources are also progressively evolved to introduce more challenging questions over time.
We highlight our contributions as follows:
  • We introduce OpenVLThinker-7B, one of the first open-source LVLMs to demonstrate reliable self-reflection, planning, and correction in visual contexts.
  • We present a simple yet effective iterative SFT-RL loop that enables R1-style reasoning into multimodal domains and steadily self-improves without requiring massive datasets.
  • We analyse linguistic markers of complex reasoning and show that SFT can steer RL exploration toward highlighted reasoning actions.
  • On six challenging benchmarks, including MathVista and MathVerse, OpenVLThinker presents remarkable improvements while reducing hallucination on HallusionBench.
@@ 2.1 Complex Chain-of-Thought Reasoning
Since the introduction of OpenAI’s O1 model [29], researchers have shown strong interest in reproducing and enhancing the complex reasoning capabilities of LLMs [51, 27, 46, 82, 81], partly due to its superior performance on mathematical benchmarks. [19] introduce the open-source DeepSeek-R1 model and investigate how RL with verifiable rewards can promote advanced chain-of-thought reasoning and reflective behaviors. This development inspired a line of research focused on open-source reproduction [44, 24, 83, 40, 60] and the analysis of such complex reasoning in mathematical problem solving [77, 71, 76, 9]. In parallel, several recent studies have similarly explored the effects of test-time scaling on encouraging more complex model reasoning behaviors [48, 53, 38, 16, 85, 58]. However, the majority of research have significantly advanced text-based reasoning, and development of vision-language reasoning is much more initial.
@@ 2.2 Vision-Language Reasoning Model
Recent advancements in large vision-language models (LVLMs) stem from open-source LLMs [63, 64, 13, 73] and text-aligned image encoders [52, 34]. Integrating these components has enabled LVLMs to follow diverse visual instructions and generate meaningful responses [36, 14, 15, 10, 37, 3]. Parallel to the model development, researchers have also been interested in eliciting CoT reasoning chains from LVLMs via prompting [87, 88, 47, 25] or fine-tuning [20, 72, 62, 12]. These reasoning models remain mostly on a shallow level of common step-by-step prompting, without self-reflections or self-verifications.
Concurrent work. Very recently, many studies have started exploring how to equip LVLMs with R1-like reasoning capabilities through distillation from text-only reasoning models [5, 74, 26, 32] or directly rely on RL [89, 42] for self-exploration. Further advancements [55, 69, 6, 45, 70, 66, 39, 84] have focused on improving performance in visual math reasoning, which marks the transition from early-stage exploration to more effective complex vision-language reasoning. Please note that most of these works are within the two months before the submission date, and some of them do not even have associated technical reports available yet. Our work aligns with these studies and contributes unique insights into the role of SFT for complex reasoning, along with an iterative SFT-RL framework to further advance research in this direction.
Figure 3: Iterative performance improvement of our model on MathVista. We note that Iter(i) is always fine-tuned from the base model Qwen2.5-VL-7B, with its training data generated from Iter(i-1).
Figure 4: An example of OpenVLThinker-7B reasoning. Question: In the diagram of the food web shown, if the number of ferns decrease, the supply of salmon will most likely? (A) decrease (B) can’t tell (C) stay same (D) increase. Corresponding image is shown in Figure 12.
@@ 3 Preliminaries
Supervised Fine-Tuning (SFT). SFT is typically applied to specialize LLMs for a particular task or domain. This process updates the model parameters $\bm{\theta}$ by providing example responses of desired behavior to the input instructions. Concretely, Given a dataset $\mathcal{D}=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N}$, where $\mathbf{x}^{(i)}$ is the prompt sequence and $\mathbf{y}^{(i)}$ is the desired response sequence. We update $\bm{\theta}$ to maximize the likelihood of producing $\mathbf{y}^{(i)}$ given $\mathbf{x}^{(i)}$. Formally, $\mathcal{L}_{\text{SFT}}(\bm{\theta})=-\sum_{i=1}^{N}\log p_{\bm{\theta}}\bigl{(}\mathbf{y}^{(i)}\,\big{|}\,\mathbf{x}^{(i)}\bigr{)}$. By minimizing the loss, the model learns to produce responses more aligned with the labeled examples.
Reinforcement Learning (RL). RL approaches fine-tune LLMs via human preferences modeled under the Bradley-Terry model [49, 11, 54, 1]: $p(\mathbf{y}_{w}\succ\mathbf{y}_{l}\mid\mathbf{x})=\sigma\bigl{(}r(\mathbf{x},\mathbf{y}_{w})-r(\mathbf{x},\mathbf{y}_{l})\bigr{)}$, where $\mathbf{y}_{w}$ and $\mathbf{y}_{l}$ denote preferred and dispreferred responses, respectively, and $\sigma(t)=1/(1+e^{-t})$ is the sigmoid function. The common RL objective under the Bradley-Terry assumption of the reward model $r(\mathbf{x},\mathbf{y})$ is thus
@@
where $\beta>0$ is the KL penalty coefficient. Under this framework, [54] introduced Group Relative Policy Optimization (GRPO) by sampling a group of response trajectories $\{\mathbf{o}_{i}\}_{i=1}^{G}$ from the old policy model $\bm{\theta}_{\mathrm{old}}$ for each query $\mathbf{x}$, with the objective as maximizing: